OpenStudy (anonymous):

Compute the nth derivative f^{(n)}(t) of the function f(t)=frac{(e^t-1)}{t}. Then compute \[\lim_{t \rightarrow 0}f^{(n)}(t)\]. This is from probability theory dealing with moment generating functions. I'm having trouble solving this.

OpenStudy (anonymous):

yuck the latex came out badly

OpenStudy (anonymous):

let me re-write

OpenStudy (anonymous):

find the nth derivative of: \[\frac{e^t-1}{t}\]

OpenStudy (anonymous):

with respect to t, and then take the limit of that nth derivative as t goes to 0

OpenStudy (anonymous):

If I was going to attempt it. I would take like 4 derivative as it stands and see if you can develop a pattern? From there, the limit will fall out.

OpenStudy (anonymous):

i took a bunch of derivatives, but there are new terms each time so I think you have to write the nth derivative as a series

OpenStudy (anonymous):

Let me try something. h/o

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hey male sory to interupt but can u check my question please...@ least where i got it sof ar

OpenStudy (anonymous):

Yeah, give me like 2 min please

OpenStudy (anonymous):

ok no prob

OpenStudy (anonymous):

purple did you post your question maybe I can help

OpenStudy (anonymous):

Well, vitale, I got half of it in terms of an nth derivative: \[\frac{d^n}{dt^n}(\frac{e^t-1}{t})=\frac{e^t(t-n)}{t^{2^n}}+(-1)^{n+1}...\]. However, I can't figure out how to handle the coefficient or powers of t on the second term. Take the first 3 derivatives. And see if what I have is right for the first part plesae.

OpenStudy (anonymous):

Look at that, I'll help purp.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i had maple compute the first 9, and I dont see your first term in the derivatives

OpenStudy (anonymous):

I might have it. Try this: \[\frac{d^n}{dt^n}(\frac{e^t-1}{t})=\frac{e^t(t-n)}{t^{2^n}}+(-1)^{n+1}(n!)t^{1(n+1)} \]. Really? At least, on through 3 these work. I broke it up and differentiated e^t/t -t^-1. Then made a pattern. Try it by hand split up. Plug in your n's and see if you get it? :/ This is the best I can do.

OpenStudy (anonymous):

If you refresh the page that won't look like pellet >.>

OpenStudy (anonymous):

let me see if it simplifies to that, one sec

OpenStudy (anonymous):

And in the last term it should be t^-(n+1) not a 1

OpenStudy (anonymous):

That works PERFECTLY for me :D

OpenStudy (anonymous):

it's not working for first derivative for me

OpenStudy (anonymous):

oh woops yes it is

OpenStudy (anonymous):

let me check the higher ones

OpenStudy (anonymous):

If you plug in 1 to my equation you get: \[\frac{e^t(t-1)}{t^2}+t^{-2}\]. Okay, I did up through 4 and they all worked. But double check it :P

OpenStudy (anonymous):

ok let me look at it for a bit

OpenStudy (anonymous):

maple is saying only the first derivative works

OpenStudy (anonymous):

the other ones are missing terms

OpenStudy (anonymous):

what did you compute as the 2nd or 3rd derivative?

OpenStudy (anonymous):

Really? Hmmm...Well for the second derivative my equation gives: \[\frac{e^t(t-2)}{t^4}-2t^{-3}\]. Which is what I get doing it by hand. It should only be 2 terms if you don't break up the fraction. Well the first derivative is: \[\frac{e^t(t-1)}{t^2}+t^{-2}\]. The derivative of that (the 2nd of the original is: \[\frac{e^t(t-1)-e^t(1)}{t^4}-2t^{-3}=\frac{e^t(t-2)}{t^4}-2t^{-3}\]. That fits doesn't it?

OpenStudy (anonymous):

let me do it by hand and check

OpenStudy (anonymous):

Okay.

myininaya (myininaya):

this is looks like it is working out nasty like lol brb

OpenStudy (anonymous):

for the second derivative I get: \[\frac{e^t(t^2-2t+2)-2}{t^3}\]

OpenStudy (anonymous):

LOL I KNOW WHY. I"M RETARDED. H/o let me retry xPPPPPP

OpenStudy (anonymous):

Sorry for the wasted time D:

OpenStudy (anonymous):

no problem im still working on it too.

OpenStudy (anonymous):

For the second I get: \[\frac{e^t(t^3-t^2-2t+1)}{t^4}\].

OpenStudy (anonymous):

this is maple's nth deriv: \[-\frac{-\left( \frac{1}{t}\right)^n (\Gamma(1+n)-\Gamma(1+n,-t))}{t}\]'

OpenStudy (anonymous):

k let me check my second deriv

OpenStudy (anonymous):

i got the same as before and just checked it on maple.

OpenStudy (anonymous):

Damn, thats an intense nth derivative. lol. GAMMA FUNCTION!!!

OpenStudy (anonymous):

yeah haha I don't know what to do with it as far as taking a limit

OpenStudy (anonymous):

I see what you mean. Limits of series's are tricky...

OpenStudy (anonymous):

well the solution to the limit should be 1/(n+1) since it's a known result in probability